OpenStudy (anonymous):

How many solutions does the absolute value equation below have? |x – 3| = 2 A. three B. two C. one D. no solution

OpenStudy (jim766):

do you know how to solve it?

OpenStudy (anonymous):

no

OpenStudy (jim766):

rewrite it as two equations x-3 = 2 or -(x-3) = 2

OpenStudy (jim766):

now solve each

OpenStudy (jim766):

x = 5 or -x+3=2 -3 -3 -x = -1 x = 1

OpenStudy (anonymous):

thank you i got it

OpenStudy (jim766):

sub each back in to check \[\left| 5-3 \right|=2\] and \[\left| 1-3 \right|=2\]

OpenStudy (jim766):

yw